OpenStudy (anonymous):

Derivative of xln^3 (x)

OpenStudy (amistre64):

just use the simple rules ... product and power

OpenStudy (amistre64):

maybe a chain for full effect

OpenStudy (anonymous):

rewrite as this... \[\sin ^{3}x=\left( \sin x \right)^{3}\] and apply the chain rule

OpenStudy (amistre64):

nah, ... id keep it in logs

OpenStudy (anonymous):

my vision sucks!!! i thought it was \[\sin ^{3}x\] sorry!

OpenStudy (anonymous):

so what is it in equation editor? \[x \ln ^{3} x \text{ ?}\] i don't think i've seen it written like that before

OpenStudy (anonymous):

\[x \left( \ln x \right)^{3}\]